If the position of an object changes with time t as x=t (t-2) then find the magnitude of net displacement of the object in the time interval t = 0 to instant when it`s velocity is zero.

Grade:

1 Answers

Eshan
5 years ago
Dear student,

x=t(t-2)=t^2-2t
\implies v=\dfrac{dx}{dt}=2t-2=0
\implies t=1s

Hence displacement from t=0s to t=1s is1^2-2(1)=-1m